### v4

# On the Lucas and Lehmer sequences in transcendental Dedekind domains

Abstract.
A Dedekind domain is said to be transcendental if it is not a field but it contains a field. In this paper, we first obtain the strong divisibility property for the Lucas and Lehmer sequences in transcendental Dedekind domains, and then establish analogues of Zsigmondy’s theorem and the primitive divisor results for such sequences in function fields.
Key words and phrases:
Primitive divisor, Zsigmondy’s theorem, Lucas sequence, Lehmer sequence, transcendental Dedekind domain, function field
2010 Mathematics Subject Classification:
11A41, 11B39, 13F05
1. Introduction
1.1. Background and motivation
Given a sequence $(a_{n})_{n\geq 1}$ of the rational integers ${\mathbb{Z}}$, it is called a divisibility sequence if $a_{m}\mid a_{n}$ whenever $m\mid n$, and it is called a strong divisibility sequence if $\gcd(a_{m},a_{n})=a_{d}$ with $d=\gcd(m,n)$ for any positive integers $m,n$. In addition, for $n\geq 2$, a prime divisor of the $n$-th term $a_{n}$ is called primitive if it does not divide $a_{k}$ for any $1\leq k<n$. The history of studying primitive prime divisors started in the 19th century with work of Bang [1] and Zsigmondy [16], which is now called Zsigmondy’s theorem and says that every term beyond the sixth in the sequence $(a^{n}-b^{n})_{n\geq 1}$ has a primitive prime divisor, where $a,b$ are positive coprime integers. This theorem was independently rediscovered by Birkhoff and Vandiver [4].
Carmichael [5] established an analogue of Zsigmondy’s theorem for the Lucas sequence $((a^{n}-b^{n})/(a-b))_{n\geq 1}$, where $a,b$ are real algebraic integers such that $a/b$ is not a root of unity, and $a+b$ and $ab$ are coprime integers in ${\mathbb{Z}}$. Ward [15] obtained such a result for the Lehmer sequence $(s_{n})_{n\geq 1}$ with $s_{n}=(a^{n}-b^{n})/(a-b)$ for odd $n$ and $s_{n}=(a^{n}-b^{n})/(a^{2}-b^{2})$ for even $n$, where $a,b$ are real, and $(a+b)^{2}$ and $ab$ are coprime integers in ${\mathbb{Z}}$. All these results, including Zsigmondy’s theorem, were extended to any number field (that is, $a$ and $b$ are relatively prime integers of a number field such that $a/b$ is not a root of unity) by Schinzel [10] in an effective but not explicit manner (see [9] for an earlier work), which was first made explicitly by Stewart [13]. Furthermore, Voutier [14, Conjecture 1] conjectured that for any integer $n>30$, the $n$-th term of a Lucas or Lehmer sequence always has a primitive prime divisor; and then in 2001, Bilu, Hanrot and Voutier [3] proved this conjecture and so listed all the Lucas and Lehmer numbers without primitive prime divisor. We refer to [2, Section 6.1] for more history about this topic.
Recently, Flatters and Ward [8] found an analogue of Zsigmondy’s theorem for a polynomial sequence $(f^{n}-g^{n})_{n\geq 1}$, where $f,g$ are two coprime polynomials in a polynomial ring $K[X]$ in one variable $X$ over any arbitrary field $K$. This was recently extended to the polynomial rings of several variables over any arbitrary field by one of us; see [11]. Moreover, the author in [11] obtained the strong divisibility property and the primitive divisor results for the Lucas and Lehmer sequences in the polynomial rings of several variables.
In this paper, we try to obtain analogues of Zsigmondy’s theorem and establish the strong divisibility property and the primitive divisor results for the Lucas and Lehmer sequences in transcendental Dedekind domains. Particularly, this includes extensions of the results in [5, 9, 10, 15] from number fields to function fields.
The approach in this paper is similar to that in [11]. However, the setting here is different (see Section 1.2), and it indeed needs some extra considerations for transcendental Dedekind domains and also for the special case of function fields.
1.2. Set-up
A Dedekind domain $R$ is said to be transcendental if it is not a field but it contains a field. Recall that $R$ is integrally closed. This implies that the fraction field of $R$ is transcendental over the largest field contained in $R$. This concept seems not appear in the literature, here we use it for simplicity.
Here we give a typical example of transcendental Dedekind domains. Let $K(X)$ be the fraction field of $K[X]$ defined above. A finite extension of $K(X)$ is called a function field. It is well-known that the integral closure of $K[X]$ in a function field (over $K(X)$) is a Dedekind domain. For simplicity, we call it the integral closure of the relevant function field. Clearly, it is transcendental. So, the main results in this paper hold automatically for function fields.
We refer to [6, Lemma 1-1 and Lemma 2-1] for more examples of transcendental Dedekind domains.
From now on, let $R$ be a transcendental Dedekind domain unless specified otherwise. Let $p$ be the characteristic of $R$. Notice that either $p=0$ or $p$ is a prime number.
An element $u\in R$ is called a unit of $R$ if $1/u\in R$. Two elements of $R$ are called coprime if they are not both divisible by the same non-unit in $R$. Two elements $a$ and $b$ in $R$ are called associated if $a=ub$ for a unit $u$ of $R$, which is equivalent to $\langle a\rangle=\langle b\rangle$. Here $\langle a\rangle$ means the ideal of $R$ generated by $a$. In addition, two ideals $I$ and $J$ of $R$ are called coprime if $I+J=R$.
As usual, a common divisor of two elements $a$ and $b$ of $R$ is an element of $R$ dividing both $a$ and $b$. A greatest common divisor of $a$ and $b$, denoted by $\gcd(a,b)$, is defined as a common divisor of $a$ and $b$ which is divisible by any other common divisor. We remark that a greatest common divisor may not exist and may be not unique. Moreover, if $a$ and $b$ have more than one greatest common divisors, then all their greatest common divisors are associated.
By the way, for any two positive integers $m$ and $n$ we still use $\gcd(m,n)$ to denote their greatest common divisor, which is a positive integer.
For any ideal $I$ and any element $b$ of $R$, we say that $I$ divides $b$ (or write $I\mid b$) if $I$ divides the ideal $\langle b\rangle$ generated by $b$ in $R$, which is equivalent to $b\in I$. Besides, a prime divisor of $b$ means a prime ideal of $R$ dividing the ideal $\langle b\rangle$.
Given a sequence $(a_{n})_{n\geq 1}$ of $R$, a prime ideal $\mathfrak{q}$ of $R$ is called a primitive prime divisor of a term $a_{n}$ if $\mathfrak{q}\mid a_{n}$ but $\mathfrak{q}$ does not divide any earlier term in the sequence. Also, the sequence is called a strong divisibility sequence if for any positive integers $m,n$ with $d=\gcd(m,n)$, we have $\gcd(a_{m},a_{n})=a_{d}$ (that is, $\gcd(a_{m},a_{n})$ exists and it can be equal to $a_{d}$).
We remark that the above concept of primitive prime divisor follows from the number field case; see [9, 10].
Finally, throughout this paper we fix an algebraic closure of the fraction field of $R$. Then, the roots of a non-constant polynomial over $R$ in this algebraic closure are called algebraic elements over $R$, and the roots of a non-constant monic polynomial over $R$ are called integral elements over $R$. In particular, for any integer $n\geq 1$, the roots of the polynomial $X^{n}-1$ over $R$ are called roots of unity.
1.3. Main results
We first present an analogue of Zsigmondy’s theorem in $R$. Let $f,g$ be non-zero elements in $R$ such that the quotient $f/g$ is not a root of unity. Assume that the principal ideals $\langle f\rangle$ and $\langle g\rangle$ in $R$ are coprime ideals which are not both equal to $R$. Define the sequence $(F_{n})_{n\geq 1}$ of $R$ with respect to $f,g$ by
$F_{n}=f^{n}-g^{n},\quad n=1,2,\ldots.$
Since $f/g$ is not a root of unity, $F_{n}\neq 0$ for any integer $n\geq 1$.
Theorem 1.1.
The sequence $(F_{n})_{n\geq 1}$ is a strong divisibility sequence.
Theorem 1.2 (Zsigmondy’s theorem in $R$).
Assume that $R$ is the integral closure of a function field. If the characteristic $p>0$ and let $F^{\prime}$ be the sequence obtained from $(F_{n})_{n\geq 1}$ by deleting the terms $F_{n}$ with $p\mid n$, then each term of $F^{\prime}$ beyond the second has a primitive prime divisor. If $p=0$, then each term of $(F_{n})_{n\geq 1}$ beyond the second has a primitive prime divisor.
We remark that the sequence $F^{\prime}$ in Theorem 1.2 might no longer be a divisibility sequence. The same holds for the sequence $G^{\prime}$ in Corollary 1.3, the sequence $L^{\prime}$ in Theorem 1.5, and the sequence $U^{\prime}$ in Theorem 1.7 below.
For $f,g$ as above, define the sequence:
$G_{n}=f^{n}+g^{n},\quad n=1,2,\ldots.$
If $p\neq 2$, then $(G_{n})_{n\geq 1}\neq(F_{n})_{n\geq 1}$. Note that $F_{2n}=F_{n}G_{n}$, which implies that each primitive prime divisor of $F_{2n}$ comes from $G_{n}$. Then, the following corollary is a direct consequence of Theorem 1.2.
Corollary 1.3.
Assume that $R$ is the integral closure of a function field. If $p>0$ and let $G^{\prime}$ be the sequence obtained from $(G_{n})_{n\geq 1}$ by deleting the terms $G_{n}$ with $p\mid n$, then each term of $G^{\prime}$ beyond the second has a primitive prime divisor. If $p=0$, then each term of $(G_{n})_{n\geq 1}$ beyond the second has a primitive prime divisor.
Now, we consider the Lucas sequences. Let $\alpha,\beta$ be non-zero algebraic elements over $R$ such that the quotient $\alpha/\beta$ is not a root of unity. Assume that both $\alpha+\beta$ and $\alpha\beta$ are in $R$, and the principal ideals $\langle\alpha+\beta\rangle$ and $\langle\alpha\beta\rangle$ in $R$ are coprime ideals which are not both equal to $R$. Define the Lucas sequence $(L_{n})_{n\geq 1}$ of $R$ with respect to $\alpha,\beta$ by
$L_{n}=\frac{\alpha^{n}-\beta^{n}}{\alpha-\beta},\quad n=1,2,\ldots.$
We remark that the Lucas sequence $(L_{n})_{n\geq 1}$ satisfies the following recurrence relation over $R$:
$L_{n+2}=(\alpha+\beta)L_{n+1}-\alpha\beta L_{n},\quad n=1,2,\ldots.$
We also remark that classically the index $n$ of the Lucas sequences starts from 00. Here for simplicity and for consistency we let it start from $1$, and we do the same thing for the Lehmer sequences below.
Theorem 1.4.
The sequence $(L_{n})_{n\geq 1}$ is a strong divisibility sequence.
Theorem 1.5.
Assume that $R$ is the integral closure of a function field. If the characteristic $p>0$ and let $L^{\prime}$ be the sequence obtained from $(L_{n})_{n\geq 1}$ by deleting the terms $L_{n}$ with $p\mid n$, then each term of $L^{\prime}$ beyond the second has a primitive prime divisor. If $p=0$, then each term of $(L_{n})_{n\geq 1}$ beyond the second has a primitive prime divisor.
Now, we consider the Lehmer sequences. Let $\lambda,\eta$ be non-zero algebraic elements over $R$ such that $\lambda/\eta$ is not a root of unity. Assume that both $(\lambda+\eta)^{2}$ and $\lambda\eta$ are in $R$, and the principal ideals $\langle(\lambda+\eta)^{2}\rangle$ and $\langle\lambda\eta\rangle$ of $R$ are coprime ideals which are not both equal to $R$. Define the Lehmer sequence $(U_{n})_{n\geq 1}$ of $R$ with respect to $\lambda,\eta$ by
$U_{n}=\left\{\begin{array}[]{ll}\dfrac{\lambda^{n}-\eta^{n}}{\lambda-\eta}&\textrm{if $n$ is odd,}\\ \\ \dfrac{\lambda^{n}-\eta^{n}}{\lambda^{2}-\eta^{2}}&\textrm{if $n$ is even.}\end{array}\right.$
We remark that the Lehmer sequence $(U_{n})_{n\geq 1}$ satisfies the following recurrence relation over $R$:
$U_{n+4}=(\lambda^{2}+\eta^{2})U_{n+2}-(\lambda\eta)^{2}U_{n},\quad n=1,2,\ldots.$
The strong divisibility property and the primitive divisor property of the sequence $(U_{n})_{n\geq 1}$ can be obtained by a similar way as in [11].
Theorem 1.6.
The sequence $(U_{n})_{n\geq 1}$ is a strong divisibility sequence.
Theorem 1.7.
Assume that $R$ is the integral closure of a function field. If the characteristic $p>0$ and let $U^{\prime}$ be the sequence obtained from $(U_{n})_{n\geq 1}$ by deleting the terms $U_{n}$ with $p\mid n$, then each term of $U^{\prime}$ beyond the second has a primitive prime divisor. If $p=0$, then each term of $(U_{n})_{n\geq 1}$ beyond the second has a primitive prime divisor.
Here we explain the reason in Theorem 1.7 that why we remove those terms with indices divisible by the characteristic $p$ when $p>0$. First, for any even integer $m\geq 2$, we have
$U_{mp}=\frac{\lambda^{mp}-\eta^{mp}}{\lambda^{2}-\eta^{2}}$ $=\frac{(\lambda^{m}-\eta^{m})^{p}}{\lambda^{2}-\eta^{2}}=\Big{(}\frac{\lambda^{m}-\eta^{m}}{\lambda^{2}-\eta^{2}}\Big{)}^{p}\cdot\frac{(\lambda^{2}-\eta^{2})^{p}}{\lambda^{2}-\eta^{2}}$
$=U_{m}^{p}\cdot\frac{\lambda^{2p}-\eta^{2p}}{\lambda^{2}-\eta^{2}}=U_{m}^{p}U_{2p}.$
Similarly, for any odd integer $m\geq 1$, we have
$U_{mp}=\left\{\begin{array}[]{ll}U_{m}^{p}U_{p}&\quad\text{if $p$ is odd},\\ U_{m}^{p}&\quad\text{if $p=2$}.\end{array}\right.$
Hence, for any integer $m>2$, the term $U_{mp}$ does not have any primitive prime divisor in the sequence $(U_{n})_{n\geq 1}$ when $p>0$.
For the proofs, we first prove Theorems 1.6 and 1.7 in Section 3, and then prove Theorems 1.4 and 1.5 in Section 4, and finally prove Theorems 1.1 and 1.2 in Section 5. We collect some basic results in Section 2.
2. Preliminaries
In this section, we collect some basic results for the proofs of the main results. Recall that $R$ is a transcendental Dedekind domain.
We first recall some basic properties of the resultant of two homogeneous polynomials in the following lemma; see [12, Proposition 2.13].
Lemma 2.1.
Let
$A(X,Y)=a_{0}X^{m}+a_{1}X^{m-1}Y+\cdots+a_{m}Y^{m}=a_{0}\prod_{i=1}^{m}(X-\alpha_{i}Y)$
and
$B(X,Y)=b_{0}X^{n}+b_{1}X^{n-1}Y+\cdots+b_{n}Y^{n}=b_{0}\prod_{j=1}^{n}(X-\beta_{j}Y)$
be two non-constant homogeneous polynomials of degrees $m$ and $n$ defined over a field. Then, their resultant $\mathrm{Res}(A,B)$ equals to
$\mathrm{Res}(A,B)=a_{0}^{n}b_{0}^{m}\prod_{i=1}^{m}\prod_{j=1}^{n}(\alpha_{i}-\beta_{j})\in{\mathbb{Z}}[a_{0},\ldots,a_{m},b_{0},\ldots,b_{n}].$
Moreover, there exist polynomials $C_{1},D_{1},C_{2},D_{2}\in{\mathbb{Z}}[a_{0},\ldots,a_{m},b_{0},\ldots,b_{n}][X,Y]$ homogeneous in $X$ and $Y$ such that
$C_{1}A+D_{1}B=\mathrm{Res}(A,B)X^{m+n-1},$
$C_{2}A+D_{2}B=\mathrm{Res}(A,B)Y^{m+n-1}.$
For any integer $n\geq 1$, the $n$-th homogeneous cyclotomic polynomial is defined by
$\Phi_{n}(X,Y)=\prod_{k=1,\,\gcd(k,n)=1}^{n}(X-\zeta_{n}^{k}Y)\in{\mathbb{Z}}[X,Y],$
where $\zeta_{n}$ is a primitive $n$-th root of unity, and we also define the polynomial
$P_{n}(X,Y)=\frac{X^{n}-Y^{n}}{X-Y}=\sum_{k=0}^{n-1}X^{n-1-k}Y^{k}=\prod_{k=1}^{n-1}(X-\zeta_{n}^{k}Y).$
Clearly,
$X^{n}-Y^{n}=\prod_{d\mid n}\Phi_{d}(X,Y),\qquad P_{n}(X,Y)=\prod_{d\mid n,\,d\geq 2}\Phi_{d}(X,Y).$
It is well-known that the degree of $\Phi(X,Y)$ equals to $\varphi(n)$, where $\varphi$ is Euler’s totient function.
The following lemma is [8, Lemma 2.4] about the resultant of $P_{m}(X,Y)$ and $P_{n}(X,Y)$.
Lemma 2.2.
For any positive coprime integers $m$ and $n$, we have $\mathrm{Res}(P_{m},P_{n})=\pm 1$.
The next lemma is a strong analogue of [11, Lemma 2.3]. It is about coprime ideals (instead of coprime elements). Note that coprime principal ideals imply coprime elements, but the converse is not true in general.
Lemma 2.3.
Let $a,b$ be two algebraic elements over $R$. Assume that both $(a+b)^{2}$ and $ab$ are in $R$, and the two principal ideals $\langle(a+b)^{2}\rangle$ and $\langle ab\rangle$ of $R$ are coprime. Let $A(X,Y),B(X,Y)\in{\mathbb{Z}}[X,Y]$ be non-constant homogeneous polynomials with resultant $\mathrm{Res}(A,B)=\pm 1$. Assume that both $A(a,b)$ and $B(a,b)$ are in $R$. Then, the principal ideals $\langle A(a,b)\rangle$ and $\langle B(a,b)\rangle$ of $R$ are coprime.
Proof.
Let $m=\deg A$ and $n=\deg B$. Note that both $(a+b)^{2}$ and $ab$ are in $R$. So, by induction we obtain that for any integer $k\geq 1$, $a^{2k}+b^{2k}$ is also in $R$. By Lemma 2.1 and noticing $\mathrm{Res}(A,B)=\pm 1$, we have that there exist $u_{1},w_{1},u_{2},w_{2}\in{\mathbb{Z}}[a,b]$ such that
$u_{1}A(a,b)+w_{1}B(a,b)=a^{2(m+n-1)}+b^{2(m+n-1)}\in R$
and
$u_{2}A(a,b)+w_{2}B(a,b)=a^{2(m+n-1)}b^{2(m+n-1)}\in R.$
Notice that $u_{1},w_{1},u_{2},w_{2}$ might be not in $R$.
By contradiction, suppose that the principal ideals $\langle A(a,b)\rangle$ and $\langle B(a,b)\rangle$ of $R$ are not coprime. Then, there is a prime ideal, say $\mathfrak{q}$, of $R$ such that both $A(a,b)$ and $B(a,b)$ are in $\mathfrak{q}$.
Now, let $M$ be the fractional field of $R$, let $M(a,b)$ be the field generated by $a$ and $b$ over $M$, and let $\overline{R}$ be the integral closure of $R$ in $M(a,b)$. Since $a,b$ are two algebraic elements over $R$, we have that $M(a,b)$ is a finite extension of $M$. So, $\overline{R}$ is also a Dedekind domain. Picking up a prime ideal, say $\overline{\mathfrak{q}}$, of $\overline{R}$, which contains $\mathfrak{q}$, we directly have that both $A(a,b)$ and $B(a,b)$ are in $\overline{\mathfrak{q}}$.
By assumption, we note that $a,b$ are integral over $R$ satisfying the equation $X^{4}-(a^{2}+b^{2})X^{2}+a^{2}b^{2}=0$. So, both $a$ and $b$ are in $\overline{R}$. Then, the above $u_{1},w_{1},u_{2},w_{2}$ are all in $\overline{R}$. Hence, we get that $a^{2(m+n-1)}+b^{2(m+n-1)}\in\overline{\mathfrak{q}}$ and $a^{2(m+n-1)}b^{2(m+n-1)}\in\overline{\mathfrak{q}}$. Moreover, since both $a^{2(m+n-1)}+b^{2(m+n-1)}$ and $a^{2(m+n-1)}b^{2(m+n-1)}$ (which equals to $(ab)^{2(m+n-1)}$) are in $R$, we know that they are both in $\mathfrak{q}$ (because $\overline{\mathfrak{q}}\cap R=\mathfrak{q}$). So, we have $ab\in\mathfrak{q}$.
Let $k=m+n-1$. We have shown that $\mathfrak{q}\mid a^{2k}+b^{2k}$ and $\mathfrak{q}\mid ab$ in $R$. Note that
$(a^{2}+b^{2})^{k}$ $=a^{2k}+b^{2k}+\sum_{i=1}^{k-1}\binom{k}{i}(a^{2})^{i}(b^{2})^{k-i}$
$=a^{2k}+b^{2k}+a^{2}b^{2}\sum_{i=1}^{k-1}\binom{k}{i}(a^{2})^{i-1}(b^{2})^{k-1-i},$
where $\sum_{i=1}^{k-1}\binom{k}{i}(a^{2})^{i-1}(b^{2})^{k-1-i}$ is in $R$ (because it is symmetric in $a^{2}$ and $b^{2}$, and both $a^{2}+b^{2}$ and $a^{2}b^{2}$ are in $R$). Hence, we obtain $\mathfrak{q}\mid(a^{2}+b^{2})^{k}$, then $\mathfrak{q}\mid a^{2}+b^{2}$, and thus $\mathfrak{q}\mid(a+b)^{2}$ in $R$ (because $\mathfrak{q}\mid ab$). This leads to a contradiction with the assumption that the principal ideals $\langle(a+b)^{2}\rangle$ and $\langle ab\rangle$ of $R$ are coprime. Therefore, the principal ideals $\langle A(a,b)\rangle$ and $\langle B(a,b)\rangle$ of $R$ are coprime. ∎
Using Lemma 2.3, we obtain the following five lemmas about coprime ideals in $R$ (instead of coprime elements in $R$), which are strong analogues of [11, Lemmas 2.4–2.8]. For the convenience of the readers, we also present their proofs.
Lemma 2.4.
Let $a,b$ be two algebraic elements over $R$. Assume that both $a+b$ and $ab$ are in $R$, and the two principal ideals $\langle a+b\rangle$ and $\langle ab\rangle$ of $R$ are coprime. Then, for any positive coprime integers $m,n$, the principal ideals $\langle P_{m}(a,b)\rangle$ and $\langle P_{n}(a,b)\rangle$ of $R$ are coprime.
Proof.
Without loss of generality, we can assume $m\geq 2,n\geq 2$. Clearly, both $P_{m}(a,b)$ and $P_{n}(a,b)$ are in $R$. Because both $a+b$ and $ab$ are in $R$, and both $P_{m}(X,Y)$ and $P_{n}(X,Y)$ are symmetric in $X$ and $Y$.
By assumption, the integers $m$ and $n$ are coprime, and the principal ideals $\langle(a+b)^{2}\rangle$ and $\langle ab\rangle$ of $R$ are coprime. Then, the desired result follows directly from Lemmas 2.2 and 2.3. ∎
Lemma 2.5.
Let $a,b$ be defined as in Lemma 2.3. Let $m,n$ be two positive coprime integers such that both $m$ and $n$ are odd. Then, the principal ideals $\langle P_{m}(a,b)\rangle$ and $\langle P_{n}(a,b)\rangle$ of $R$ are coprime.
Proof.
Without loss of generality, we can assume $m\geq 3,n\geq 3$. By assumption, both $a^{2}+b^{2}$ and $ab$ are in $R$. Moreover, we have that for any integer $k\geq 1$, $a^{2k}+b^{2k}\in R$.
Note that $P_{m}(X,Y)$ is homogeneous of even degree $m-1$ (since $m$ is odd) and symmetric in $X$ and $Y$. So, if $X^{i}Y^{j}$ is a term in $P_{m}(X,Y)$, then $X^{j}Y^{i}$ is also a term in $P_{m}(X,Y)$, and then assuming $i\leq j$, we have
$a^{i}b^{j}+a^{j}b^{i}=(ab)^{i}(a^{j-i}+b^{j-i})\in R,$
where we use the fact that $j-i$ is even (because $i+j=m-1$ is even). Hence, we have that $P_{m}(a,b)$ is in $R$. Similarly, $P_{n}(a,b)$ is also in $R$, because $n$ is also odd.
Now, the desired result follows directly from Lemmas 2.2 and 2.3. ∎
Lemma 2.6.
Let $a,b$ be defined as in Lemma 2.3. Let $m,n$ be two positive coprime integers such that $m$ is odd and $n$ is even. Then, the principal ideals $\langle P_{m}(a,b)\rangle$ and $\langle P_{n}(a,b)/(a+b)\rangle$ of $R$ are coprime.
Proof.
Without loss of generality, we can assume $m\geq 3,n\geq 2$. Since $m$ is odd, as in the proof of Lemma 2.5 $P_{m}(a,b)$ is in $R$. For any odd integer $k\geq 1$, note that
$\frac{a^{k}+b^{k}}{a+b}=a^{k-1}-a^{k-2}b+\cdots-ab^{k-2}+b^{k-1}$
is homogeneous of even degree $k-1$ and is symmetric in $a$ and $b$, so it is in $R$. Hence, for even $n$, since
$\frac{P_{n}(a,b)}{a+b}=\frac{a^{n-1}+b^{n-1}}{a+b}+ab\cdot\frac{a^{n-3}+b^{n-3}}{a+b}+\cdots+a^{\frac{n-2}{2}}b^{\frac{n-2}{2}}\cdot\frac{a+b}{a+b},$
we have that $P_{n}(a,b)/(a+b)$ is in $R$.
Denote $T_{n}(X,Y)=P_{n}(X,Y)/(X+Y)$, which can be viewed as a polynomial over ${\mathbb{Z}}$. It has been shown in the proof of [11, Lemma 2.6] that the resultant of $P_{m}(X,Y)$ and $T_{n}(X,Y)$ is equal to ±$\pm 1$.
Hence, using Lemma 2.3 we obtain that the principal ideals $\langle P_{m}(a,b)\rangle$ and $\langle P_{n}(a,b)/(a+b)\rangle$ of $R$ are coprime. ∎
Lemma 2.7.
Let $a,b$ be defined as in Lemma 2.3. Let $m,n$ be two positive integers such that both $m$ and $n$ are odd. Then, the principal ideals $\langle P_{m}(a^{n},b^{n})\rangle$ and $\langle(a^{n}+b^{n})/(a+b)\rangle$ of $R$ are coprime.
Proof.
Without loss of generality, we can assume $m\geq 3,n\geq 3$. As before, since $m$ and $n$ are odd, $P_{m}(a^{n},b^{n})$ and $(a^{n}+b^{n})/(a+b)$ are indeed in $R$.
Define
$V_{m}(X,Y)=P_{m}(X^{n},Y^{n}),\qquad W_{n}(X,Y)=\frac{X^{n}+Y^{n}}{X+Y}.$
It has been shown in the proof of [11, Lemma 2.7] that the resultant $\mathrm{Res}(V_{m},W_{n})=\pm 1$. So, the desired result follows directly from Lemma 2.3. ∎
Lemma 2.8.
Let $a,b$ be defined as in Lemma 2.3. Then, for any odd integer $n\geq 1$, the principal ideals $\langle P_{n}(a,b)\rangle$ and $\langle(a+b)^{2}\rangle$ of $R$ are coprime.
Proof.
Without loss of generality, we fix an odd integer $n\geq 3$. Since $n$ is odd, by Lemma 2.5 we have that $P_{n}(a,b)$ is in $R$. Moreover, it has been shown in the proof of [11, Lemma 2.8] that the resultant of the polynomials $P_{n}(X,Y)$ and $(X+Y)^{2}$ is equal to ±$\pm 1$. Hence, the desired result follows from Lemma 2.3. ∎
3. Proofs of Theorems 1.6 and 1.7
We first prove Theorem 1.6.
Proof of Theorem 1.6.
Let $d=\gcd(m,n)$.
First, assume that both $m$ and $n$ are even. Then, $d$ is also even. Clearly,
$U_{m}=U_{d}P_{m/d}(\lambda^{d},\eta^{d}),\quad U_{n}=U_{d}P_{n/d}(\lambda^{d},\eta^{d}).$
By assumption, the principal ideals $\langle\lambda^{d}+\eta^{d}\rangle$ and $\langle\lambda^{d}\eta^{d}\rangle$ of $R$ are coprime (as in the last paragraph of the proof of Lemma 2.3). Hence, it follows from Lemma 2.4 that $P_{m/d}(\lambda^{d},\eta^{d})$ and $P_{n/d}(\lambda^{d},\eta^{d})$ are coprime in $R$, and thus, $\gcd(U_{m},U_{n})=U_{d}$.
Now, assume that both $m$ and $n$ are odd. Then, $d$ is also odd. Clearly,
$U_{m}=U_{d}P_{m/d}(\lambda^{d},\eta^{d}),\quad U_{n}=U_{d}P_{n/d}(\lambda^{d},\eta^{d}).$
Note that the principal ideals $\langle(\lambda^{d}+\eta^{d})^{2}\rangle$ and $\langle\lambda^{d}\eta^{d}\rangle$ of $R$ are coprime. So, it follows from Lemma 2.5 that $P_{m/d}(\lambda^{d},\eta^{d})$ and $P_{n/d}(\lambda^{d},\eta^{d})$ are coprime in $R$, and so, $\gcd(U_{m},U_{n})=U_{d}$.
Finally, without loss of generality, we assume that $m$ is odd and $n$ is even. Then, $d$ is odd. Clearly,
$U_{m}=U_{d}P_{m/d}(\lambda^{d},\eta^{d}),$
$U_{n}=U_{d}\cdot\frac{P_{n/d}(\lambda^{d},\eta^{d})}{\lambda^{d}+\eta^{d}}\cdot\frac{\lambda^{d}+\eta^{d}}{\lambda+\eta}.$
Using Lemma 2.6 we have that $P_{m/d}(\lambda^{d},\eta^{d})$ and $P_{n/d}(\lambda^{d},\eta^{d})/(\lambda^{d}+\eta^{d})$ are coprime in $R$. In addition, by Lemma 2.7 we obtain that $P_{m/d}(\lambda^{d},\eta^{d})$ and $(\lambda^{d}+\eta^{d})/(\lambda+\eta)$ are coprime in $R$. Hence, $\gcd(U_{m},U_{n})=U_{d}$. This completes the proof. ∎
To prove Theorem 1.7, we need to make some more preparations.
Recall that $p$ is the characteristic of $R$. As usual, denote by $v_{\mathfrak{q}}(h)$ the maximal power to which a prime ideal $\mathfrak{q}$ divides $h\in R$.
In this section, let $M$ be the fraction field of $R$. The elements $\lambda$ and $\eta$ have been defined at the beginning of Section 1.3. By assumption, $M(\lambda)$ is an algebraic extension generated by $\lambda$ over $M$ and also containing $\eta$. For any prime ideal $\mathfrak{q}\in R$, as usual $v_{\mathfrak{q}}$ induces a non-Archimedean valuation of $M$. This valuation can be extended to the field $M(\lambda)$ (still denoted by $v_{\mathfrak{q}}$); see, for instance, [7, Theorem 3.1.2].
Lemma 3.1.
Let $\mathfrak{q}$ be a prime ideal of $R$ dividing $U_{n}$ for some $n\geq 3$. Then, for any integer $m\geq 1$ with $p\nmid m$ (normal-(((including the case $p=0)$, we have $v_{\mathfrak{q}}(U_{mn})=v_{\mathfrak{q}}(U_{n})$.
Proof.
First, note that $\lambda,\eta$ are both integral over the ring $R$ (because they satisfy the equation $X^{4}-(\lambda^{2}+\eta^{2})X^{2}+\lambda^{2}\eta^{2}=0$ over $R$). So, we have $v_{\mathfrak{q}}(\lambda)\geq 0$ and $v_{\mathfrak{q}}(\eta)\geq 0$.
Assume that $v_{\mathfrak{q}}(\eta)>0$. Note that by definition, either $\lambda^{n}=\eta^{n}+(\lambda-\eta)U_{n}$, or $\lambda^{n}=\eta^{n}+(\lambda^{2}-\eta^{2})U_{n}$. We also note that $v_{\mathfrak{q}}(U_{n})>0$ by assumption. So, we obtain $v_{\mathfrak{q}}(\lambda^{n})>0$, which gives $v_{\mathfrak{q}}(\lambda)>0$. Thus,
$v_{\mathfrak{q}}(\lambda+\eta)>0,\qquad v_{\mathfrak{q}}(\lambda\eta)>0.$
This contradicts the assumption that the principal ideals $\langle(\lambda+\eta)^{2}\rangle$ and $\langle\lambda\eta\rangle$ of $R$ are coprime. So, we must have $v_{\mathfrak{q}}(\eta)=0$.
Now, suppose that $n$ is odd. Then, $U_{n}=(\lambda^{n}-\eta^{n})/(\lambda-\eta)$, and so
$\lambda^{n}=\eta^{n}+(\lambda-\eta)U_{n}.$
Then, we have
$\lambda^{mn}=\big{(}\eta^{n}+(\lambda-\eta)U_{n}\big{)}^{m}=\eta^{mn}+\sum_{i=1}^{m}\binom{m}{i}(\lambda-\eta)^{i}U_{n}^{i}\eta^{n(m-i)},$
which implies
$\frac{\lambda^{mn}-\eta^{mn}}{\lambda-\eta}=m\eta^{n(m-1)}U_{n}+\sum_{i=2}^{m}\binom{m}{i}(\lambda-\eta)^{i-1}\eta^{n(m-i)}U_{n}^{i}.$
Hence, we have that for odd $m$
$U_{mn}=m\eta^{n(m-1)}U_{n}+\sum_{i=2}^{m}\binom{m}{i}(\lambda-\eta)^{i-1}\eta^{n(m-i)}U_{n}^{i},$
and for even $m$
$(\lambda+\eta)U_{mn}=m\eta^{n(m-1)}U_{n}+\sum_{i=2}^{m}\binom{m}{i}(\lambda-\eta)^{i-1}\eta^{n(m-i)}U_{n}^{i}.$
Besides, since $n$ is odd and $v_{\mathfrak{q}}(U_{n})>0$, using Lemma 2.8 we obtain $v_{\mathfrak{q}}(\lambda+\eta)=0$. Hence, the desired result follows from the ultrametric inequality for non-Archimedean valuations. Here we need to use $p\nmid m$ and $v_{\mathfrak{q}}(\eta)=0$.
Finally, assume that $n$ is even. Then, as the above, for any integer $m\geq 1$ we have
$U_{mn}=m\eta^{n(m-1)}U_{n}+\sum_{i=2}^{m}\binom{m}{i}(\lambda^{2}-\eta^{2})^{i-1}\eta^{n(m-i)}U_{n}^{i}.$
The desired result now follows. ∎
The following lemma is about the element $\Phi_{n}(\lambda,\eta)$.
Lemma 3.2.
For any integer $n\geq 3$, $\Phi_{n}(\lambda,\eta)$ is in $R$; and moreover, when $R$ is the integral closure of a function field, $\Phi_{n}(\lambda,\eta)$ is not a unit in $R$.
Proof.
First, note that for any integer $n\geq 3$, the polynomial $\Phi_{n}(X,Y)$ is homogeneous of even degree $\varphi(n)$ and is symmetric in $X$ and $Y$. Then, as in the proof of Lemma 2.5, we have that $\Phi_{n}(\lambda,\eta)$ is in $R$.
In the sequel, we assume that $R$ is the integral closure of a function field.
Let $K$ be the largest field contained in $R$. By the choice of $R$, we can choose a valuation at infinity, say $v_{\infty}$, of $R$ such that for any element $a\in R$, we have $v_{\infty}(a)=0$ if $a\in K$, and otherwise $v_{\infty}(a)<0$. Moreover, for any integral element $b$ over $R$, we have $v_{\infty}(b)<0$ if $b$ is not algebraic over $K$.
By assumption, we know that the quotient $\lambda/\eta$ is not a root of unity, and the principal ideals $\langle(\lambda+\eta)^{2}\rangle$ and $\langle\lambda\eta\rangle$ of $R$ are coprime ideals which are not both equal to $R$. Then, $(\lambda+\eta)^{2}$ and $\lambda\eta$ can not be both in $K$. So, at least one of $\lambda$ and $\eta$ is transcendental over $K$.
Without loss of generality, we assume that $\lambda$ is transcendental over $K$. Note that $\lambda$ is integral over $R$. Then, we have
$v_{\infty}(\lambda)<0.$
Recall that $M$ is the fractional field of $R$, and $M(\zeta_{n},\lambda)$ is an algebraic extension of $M$ containing $\eta$. Then, it is well-known that the valuation $v_{\infty}$ of $R$ can be extended to $M(\zeta_{n},\lambda)$ (see, for instance, [7, Theorem 3.1.2]), and we still denote this extension by $v_{\infty}$.
By definition, we have
$\Phi_{n}(\lambda,\eta)=\prod_{\begin{subarray}{c}k=1\\ \gcd(k,n)=1\end{subarray}}^{n}(\lambda-\zeta_{n}^{k}\eta).$
Then, we have
$v_{\infty}(\Phi_{n}(\lambda,\eta))=\sum_{\begin{subarray}{c}k=1\\ \gcd(k,n)=1\end{subarray}}^{n}v_{\infty}(\lambda-\zeta_{n}^{k}\eta).$
If $v_{\infty}(\lambda)\neq v_{\infty}(\eta)$, then we obtain that for any integer $k\geq 1$,
$v_{\infty}(\lambda-\zeta_{n}^{k}\eta)=\min\{v_{\infty}(\lambda),v_{\infty}(\eta)\}\leq v_{\infty}(\lambda)<0,$
which implies
$v_{\infty}(\Phi_{n}(\lambda,\eta))<0.$
Now, assume that $v_{\infty}(\lambda)=v_{\infty}(\eta)=m$ (note that $m<0$). First, we note that $\lambda$ and $\eta$ are both integral over $R$, and so is $\lambda-\zeta_{n}^{k}\eta$ for any integer $k\geq 1$. So, we have that for any integer $k\geq 1$,
(3.1) $v_{\infty}(\lambda-\zeta_{n}^{k}\eta)\leq 0.$
Since $n\geq 3$, we have $\varphi(n)\geq 2$, and so $\Phi_{n}(\lambda,\eta)$ has at least two distinct factors, say $\lambda-\zeta_{n}^{k_{1}}\eta$ and $\lambda-\zeta_{n}^{k_{2}}\eta$ with $\zeta_{n}^{k_{1}}\neq\zeta_{n}^{k_{2}}$. Using the ultrametric inequality for non-Archimedean valuations, we have
$v_{\infty}(\lambda-\zeta_{n}^{k_{1}}\eta)\geq\min\{v_{\infty}(\lambda),v_{\infty}(\eta)\}=m.$
If $v_{\infty}(\lambda-\zeta_{n}^{k_{1}}\eta)=m$, then noticing (3.1) we get $v_{\infty}(\Phi_{n}(\lambda,\eta))\leq m<0$. If $v_{\infty}(\lambda-\zeta_{n}^{k_{1}}\eta)>m$, then noticing $\lambda-\zeta_{n}^{k_{2}}\eta=(\lambda-\zeta_{n}^{k_{1}}\eta)+(\zeta_{n}^{k_{1}}-\zeta_{n}^{k_{2}})\eta$, we obtain
$v_{\infty}(\lambda-\zeta_{n}^{k_{2}}\eta)=v_{\infty}((\zeta_{n}^{k_{1}}-\zeta_{n}^{k_{2}})\eta)=v_{\infty}(\eta)=m,$
which, together with (3.1), gives $v_{\infty}(\Phi_{n}(\lambda,\eta))\leq m<0$.
Therefore, we always have $v_{\infty}(\Phi_{n}(\lambda,\eta))<0$. Then, we must have that $\Phi_{n}(\lambda,\eta)$ is not a unit in $R$. Because any unit of $R$ has zero value in any valuation of $R$. ∎
Remark 3.3.
The proof of the second part of Lemma 3.2 relies on the property of a valuation at infinity $v_{\infty}$ of $R$ that $v_{\infty}(a)\leq 0$ for any $a\in R$. This property does not hold in general for transcendental Dedekind domains. For example, let $K[X]$ be a polynomial ring in one variable $X$ over a field $K$, let $S=\{h^{k}:\,k\in{\mathbb{Z}},k\geq 0\}$ where $h$ is a fixed irreducible polynomial in $K[X]$, and let $R$ be the localization of $K[X]$ in the set $S$, then by [6, Lemma 1-1] we know that $R$ is a Dedekind domain, and clearly it is a transcendental Dedekind domain, but for the unique valuation at infinity $v_{\infty}$ of $R$ corresponding to $1/X$, we have $v_{\infty}(1/h)=\deg h>0$, where $1/h\in R$.
The following lemma asserts that when the characteristic $p\nmid n$ (including the case $p=0$), any non-primitive prime divisor of the term $U_{n}$ does not divide $\Phi_{n}(\lambda,\eta)$. Recall that when $p>0$, $U^{\prime}$ is the sequence obtained from $(U_{n})_{n\geq 1}$ by deleting the terms $U_{n}$ with $p\mid n$,
Lemma 3.4.
Let $n$ be an integer such that $n\geq 3$ and $p\nmid n$ if $p>0$. If $p>0$, then any non-primitive prime divisor of the term $U_{n}$ (normal-(((in the sequence $U^{\prime})$ does not divide $\Phi_{n}(\lambda,\eta)$. If $p=0$, then any non-primitive prime divisor of the term $U_{n}$ does not divide $\Phi_{n}(\lambda,\eta)$.
Proof.
As in [15], we define the sequence $(Q_{k})_{k\geq 1}$ of polynomials by $Q_{1}=1,Q_{2}=1$, and
$Q_{k}(X,Y)=\Phi_{k}(X,Y),\quad k=3,4,\ldots.$
Clearly, for integer $n\geq 3$ we have
$U_{n}=\prod_{d\mid n}Q_{d}(\lambda,\eta).$
By the Möbius inversion we obtain
$Q_{n}(\lambda,\eta)=\prod_{d\mid n}U_{d}^{\mu(n/d)},$
where $\mu$ is the Möbius function.
Now, suppose that the characteristic $p>0$. Let $\mathfrak{q}$ be a prime divisor of $U_{n}$ which is not primitive (in the sequence $U^{\prime}$), where $p\nmid n$. Let $m$ be the minimal positive integer such that $\mathfrak{q}\mid U_{m}$ and $p\nmid m$. Since $\mathfrak{q}$ is not primitive, we have $m<n$. Moreover, using Theorem 1.6 we obtain that $\gcd(U_{m},U_{n})=U_{\gcd(m,n)}$, and so $m\mid n$ (noticing the minimality of $m$). Similarly, for any positive integer $d$, we obtain $v_{\mathfrak{q}}(U_{d})=0$ if $m\nmid d$. Besides, by Lemma 3.1, for any $d$ with $p\nmid d$ we have
$v_{\mathfrak{q}}(U_{dm})=v_{\mathfrak{q}}(U_{m}).$
Hence, noticing $p\nmid n$ we obtain
$v_{\mathfrak{q}}(Q_{n}(\lambda,\eta))$ $=\sum_{d\mid n}\mu(n/d)v_{\mathfrak{q}}(U_{d})$
$=\sum_{dm\mid n}\mu(n/(dm))v_{\mathfrak{q}}(U_{dm})$
$=\sum_{d\mid n/m}\mu(n/(dm))v_{\mathfrak{q}}(U_{m})$
$=v_{\mathfrak{q}}(U_{m})\sum_{d\mid n/m}\mu(n/(dm))=0.$
So, any non-primitive prime divisor of $U_{n}$ (in the sequence $U^{\prime}$) does not divide $Q_{n}(\lambda,\eta)$ (that is, $\Phi_{n}(\lambda,\eta)$, since $n\geq 3$).
The proof for the case $p=0$ follows exactly the same way. ∎
Now, we are ready to prove Theorem 1.7.
Proof of Theorem 1.7.
Recall that $Q_{1}=1,Q_{2}=1$, and $Q_{k}(X,Y)=\Phi_{k}(X,Y),\quad k=3,4,\ldots$. Note that for integer $n\geq 3$ we have
$U_{n}=\prod_{d\mid n}Q_{d}(\lambda,\eta).$
So, any primitive prime divisor of the term $U_{n}$ (in the sequence $U^{\prime}$ if $p>0$) must divide $Q_{n}(\lambda,\eta)$ (that is, $\Phi_{n}(\lambda,\eta)$).
In addition, by Lemma 3.4 we know that any non-primitive prime divisor of $U_{n}$ (in the sequence $U^{\prime}$ if $p>0$) does not divide $\Phi_{n}(\lambda,\eta)$.
Hence, the primitive prime divisors of $U_{n}$ (in the sequence $U^{\prime}$ if $p>0$) are exactly the prime divisors of $\Phi_{n}(\lambda,\eta)$ (if they exist).
Note that here $R$ is the integral closure of a function field. By Lemma 3.2 and noticing $n\geq 3$, we know that $\Phi_{n}(\lambda,\eta)$ is in $R$ and it is not a unit in $R$. So, $\Phi_{n}(\lambda,\eta)$ has a prime divisor in $R$.
Therefore, collecting the above deductions, we conclude that each term beyond the second (in the sequence $U^{\prime}$ if $p>0$) has a primitive prime divisor. ∎
Remark 3.5.
In the proof of Theorem 1.7, we obtain more: when $R$ is the integral closure of a function field, the primitive part (that is, the product of all the primitive prime divisors to their respective powers) of $U_{n}$ is exactly the principal ideal $\langle\Phi_{n}(\lambda,\eta)\rangle$, where $n\geq 3$, and $p\nmid n$ if $p>0$.
4. Proofs of Theorems 1.4 and 1.5
Both proofs follow from the same arguments as in [11, Section 4] with only minor changes and from Theorems 1.6 and 1.7.
Proof of Theorem 1.4.
Fix positive integers $m,n$ with $d=\gcd(m,n)$. If either both $m,n$ are even, or both $m,n$ are odd, it follows directly from Theorem 1.6 that $\gcd(L_{m},L_{n})=L_{d}$ (choosing $\lambda=\alpha,\eta=\beta$).
Now, without loss of generality, assume that $m$ is even and $n$ is odd. Then, $d$ is odd. By Theorem 1.6 we obtain
$\gcd\Big{(}\frac{\alpha^{m}-\beta^{m}}{\alpha^{2}-\beta^{2}},\frac{\alpha^{n}-\beta^{n}}{\alpha-\beta}\Big{)}=\frac{\alpha^{d}-\beta^{d}}{\alpha-\beta}.$
Using Lemma 2.8 we know that $(\alpha^{n}-\beta^{n})/(\alpha-\beta)$ and $\alpha+\beta$ are coprime in $R$. Hence, we obtain $\gcd(L_{m},L_{n})=L_{d}$. This completes the proof. ∎
Proof of Theorem 1.5.
Assume that the characteristic $p=0$. First, by Lemma 3.2 (choosing $\lambda=\alpha,\eta=\beta$ there), we know that for any integer $n\geq 3$, $\Phi_{n}(\alpha,\beta)$ is in $R$ and it is not a unit in $R$.
For any odd integer $n\geq 3$, by Lemma 2.8 and noticing $\Phi_{n}(\alpha,\beta)\mid P_{n}(\alpha,\beta)$, we have that the principal ideals $\langle\Phi_{n}(\alpha,\beta)\rangle$ and $\langle\alpha+\beta\rangle$ of $R$ are coprime.
Now, fix an even integer $n\geq 4$. Suppose that there exists a prime ideal $\mathfrak{q}$ of $R$ dividing both $\Phi_{n}(\alpha,\beta)$ and $\Phi_{2}(\alpha,\beta)=\alpha+\beta$. This means that the polynomial $X^{n}-Y^{n}=\prod_{d\mid n}\Phi_{d}(X,Y)$, viewed it as a polynomial over the quotient field $R/\mathfrak{q}R$, has a multiple root (that is, $(\alpha,\beta)$). However, this field has characteristic zero (because $R$ contains a field whose characteristic is zero), which implies that $X^{n}-Y^{n}$ is in fact a simple polynomial. Hence, this leads to a contradiction, and so $\langle\Phi_{n}(\alpha,\beta)\rangle$ and $\langle\alpha+\beta\rangle$ are coprime ideals in $R$.
Hence, we always have that for any integer $n\geq 3$, the ideals $\langle\Phi_{n}(\alpha,\beta)\rangle$ and $\langle\alpha+\beta\rangle$ are coprime in $R$.
Therefore, by construction we directly obtain from Remark 3.5 that the primitive part of $L_{n}$ is exactly the ideal $\langle\Phi_{n}(\alpha,\beta)\rangle$, where $n\geq 3$.
Finally, if $p>0$, then by contruction the above arguments still work (because in the sequence $L^{\prime}$ we have deleted those terms $L_{n}$ with $p\mid n$). ∎
Remark 4.1.
From the proof, we know that when $R$ is the integral closure of a function field, the primitive part of $L_{n}$ is exactly the principal ideal $\langle\Phi_{n}(\alpha,\beta)\rangle$, where $n\geq 3$, and $p\nmid n$ if $p>0$.
5. Proofs of Theorems 1.1 and 1.2
First, Theorem 1.1 follows directly from Theorem 1.4.
For the proof of Theorem 1.2, we apply the same arguments as in [11, Proof of Theorem 1.6] and use Theorem 1.5 and Remark 4.1 (taking $\alpha=f$ and $\beta=g$).
Proof of Theorem 1.2.
Assume that the characteristic $p=0$. Fix an integer $n\geq 3$. Taking $\alpha=f$ and $\beta=g$ in Theorem 1.5 and noticing Remark 4.1, we know that the primitive part of the term $(f^{n}-g^{n})/(f-g)$ is the ideal $\langle\Phi_{n}(f,g)\rangle$. As in the proof of Theorem 1.5 and noticing $\Phi_{1}(f,g)=f-g$, we obtain that the ideals $\langle\Phi_{n}(f,g)\rangle$ and $\langle f-g\rangle$ are coprime in $R$. Hence, the primitive part of the term $F_{n}=f^{n}-g^{n}$ is the ideal $\langle\Phi_{n}(f,g)\rangle$.
Finally, if $p>0$, then by contruction the above arguments still work (because in the sequence $F^{\prime}$ we have deleted those terms $F_{n}$ with $p\mid n$). ∎
Remark 5.1.
From the proof, we know that when $R$ is the integral closure of a function field, the primitive part of $F_{n}$ is exactly the principal ideal $\langle\Phi_{n}(f,g)\rangle$, where $n\geq 3$, and $p\nmid n$ if $p>0$.
Acknowledgement
The authors would like to thank the referee for careful reading of their paper and valuable comments. For the research, Xiumei Li was supported by the National Science Foundation of China Grant No.12001312; and Min Sha was supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2022A1515012032) and also by the Australian Research Council Grant DE190100888.